OpenStudy (anonymous):

If u and v are the vectors below, find the vector w whose tail is at the point halfway from the tip of v to the tip of u and whose head is at the point halfway from the tip of u to the tip of u-v. u = (-2,-4,-5) v = (-3, 5, -1) answer: (2, -7, -3/2)

OpenStudy (anonymous):

Find two vectors pointing to those midpoints. Then subtract the vectors from each other.

OpenStudy (anonymous):

so, (1/2(U+V))-(1/2(U-V))?

OpenStudy (anonymous):

Not sure how you got that.

OpenStudy (anonymous):

ok, so one half of the way from u to v i translated into 1/2U+1/2V. The second half from u-v, based on one half of the way along U. That is 1/2U-1/2V. THen I subtracted the two from each other. Does that make sense?

OpenStudy (anonymous):

So we know \(\mathbf{w}\) is the difference between head and tail.\[ \mathbf{w}=\mathbf{w}_h-\mathbf{w}_t \]Then we know: "whose tail is at the point halfway from the tip of v to the tip of u" \[ \mathbf{w}_t=\frac{1}{2}(\mathbf{u}+\mathbf{v}) \] "whose head is at the point halfway from the tip of u to the tip of u-v"\[ \mathbf{w}_h=\frac{1}{2}(\mathbf{u}+(\mathbf{u}-\mathbf{v})) \] Does this model make sense?

OpenStudy (anonymous):

Ohhhhh. I see. I forgot to do 1/2U. And thats important because half way up U

OpenStudy (anonymous):

Wait I think my midpoint formula is wrong.

OpenStudy (anonymous):

whast wrong with it?

OpenStudy (anonymous):

whats*

OpenStudy (anonymous):

Wait never mind it is correct, false alarm.

OpenStudy (anonymous):

Ok good. I will try it and let you know

OpenStudy (anonymous):

it doesnt seem to work

OpenStudy (anonymous):

ya ive attempted the question three times now, and none of the answers I get are correct